solve for x, please show work:
1) 15x^2-8x-12=0
2) x^2-10=3x
3) 4-6(2x-8)=x-9
factor, please show work:
4) x^2-16 answer = {-4,4} I don't know if that is the correct answer
5) 3x^2-22x+7

Answer by sachi(548)

solve for x, please show work:
1) 15x^2-8x-12=0
2) x^2-10=3x
3) 4-6(2x-8)=x-9
factor, please show work:
4) x^2-16 answer = {-4,4} I don't know if that is the correct answer
5) 3x^2-22x+7
ans:-
1)15x^2-8x-12=0
or 15x^2-18x+10x-12=0
or 3x(5x-6)+2(5x-6)=0
or (5x-6)(3x+2)=0
or 5x-6=0 => x=6/5
or 3x+2=0 => x=-2/3 ans
2) x^2-10=3x
or x^2-10-3x=0
or x^2-5x+2x-10=0
or x(x-5)+2(x-5)=0
or (x-5)(x+2)=0
or x=5 or -2 ans
3) 4-6(2x-8)=x-9
or 4-12x+48-x+9=0
or -13x+61=0
or x=61/13 ans
4) x^2-16 answer = {-4,4} I don't know if that is the correct answer
yes x^2-16=0
or x^2-4^2=0
or (x+4)(x-4)=0
or x=4 or -4 [as a^2-b^2=(a+b)(a-b)] so correct
5) 3x^2-22x+7=3x^2-x-21x+7=x(3x-1)-7(3x-1)=(3x-1)(x-7) ans